# Stable and Accurate Orbital-Free DFT Powered by Machine Learning

### v1

@@ Abstract
Hohenberg and Kohn have proven that the electronic energy and the one-particle electron density can, in principle, be obtained by minimizing an energy functional with respect to the density. Given that decades of theoretical work have so far failed to produce this elusive exact energy functional promising great computational savings, it is reasonable to try and learn it empirically. Using rotationally equivariant atomistic machine learning, we obtain for the first time a density functional that, when applied to the organic molecules in QM9, yields energies with chemical accuracy while also converging to meaningful electron densities. Augmenting the training data with densities obtained from perturbed potentials proved key to these advances. Altogether, we are now closer than ever to fulfilling Hohenberg and Kohn’s promise, paving the way for more efficient calculations in large molecular systems.
@@
Sadly, the proof of existence is non-constructive. That is, we know that a universal energy functional $F[\rho]$ exists; but its form remains unknown except for simple special cases [2, 3, 4] that do not cover most chemistries of general interest.
@@
The lure of a simple but complete description of molecular ground states in terms of their electron density alone has motivated a long search for numerical approximations [7, 8]. Much past work has focused on trying to approximate energy densities in terms of the local electron density and its gradients, so-called gradient expansions. Their systematic underfitting of the data suggests the inadequacy of low-order gradient expansions [7]. The currently most successful nonempirical approach is APBEK [9] which is remarkably good given its simple form, though still far from chemical accuracy.
This motivates empirically modeling density functionals with machine learning [10, 11]: The equivariant architecture [12] was the first to demonstrate chemical accuracy of a single learned functional across input densities and geometries of tiny molecules. This work was superseded by the milestone M-OFDFT pipeline [13]. Its representation of the density in terms of a linear combination of atomic basis functions [14, 15] is much more compact, compared to a grid. It first predicted molecular energies across the QM9 dataset of diverse organic molecules with up to nine second-period atoms [16, 17] with chemical accuracy, a remarkable feat. Perhaps the biggest success of that work was its ability to extrapolate to larger molecules than trained on. One drawback was the model being fully non-local, i.e., each atom needs to exchange information with all others. While not a problem for molecules up to a few hundred atoms, this eventually becomes prohibitive for larger systems. Its foremost limitation, however, was that the learned functional did not afford variational density optimization: Following the energy gradient led to unphysical densities and energies, creating the need to engineer a method that picks an intermediate solution as the final prediction post hoc.
Figure 1: The STRUCTURES25 pipeline enables converging orbital-free density optimization. (A) Radar plot of total energy error, $L_{2}$ density error and percentage of convergence failures evaluated on QM9. STRUCTURES25 was trained on our perturbed QM9 dataset. Both energy and density error are w.r.t. the ground state labels in the orbital-free density basis (Eq. 2). (B) The STRUCTURES25 pipeline takes a molecular graph $\mathcal{M}$ and a density represented by coefficients $\mathbf{p}$ of atom-centered basis functions $\{\omega_{\mu}\}$ as input and predicts the target energy, $E_{\mathrm{TXC}}$. The gradient of the energy is obtained by automatic differentiation and used to iteratively find the ground state in density optimization.
@@
Ensuring that the learned energy functional has a true minimum at the correct ground state electron density (or very nearby) is of paramount practical importance (Fig. 2A). Indeed, it enables convergent variational density optimization to meaningful densities (Fig. 2B). If no such minimum were present, following the gradient on such a misleading surface would result in unphysical densities, e.g., allowing electrons to collapse into a nucleus. Previously, such malformed energy surfaces required elaborate procedures to salvage a prediction from a diverging density optimization trajectory [13]. In addition, an estimated ground state density with vanishing gradients $\nabla_{\mathbf{p}}E=0$ is vital for the calculation of Pulay forces, which are required for accurate geometry optimization when using atom-centered basis functions [18].
@@
yields a consistent tuple of potential $V_{\mathrm{eff}}$, orbitals $\phi_{i}$ and associated energies $\varepsilon_{i}$ from which a training sample (density coefficients $\mathbf{p}$, energy $E$, gradient $\nabla_{\mathbf{p}}E$) can be obtained. The bottom of Fig. 3C characterizes the resulting high-dimensional training data in terms of a single axis, the energy difference between training and respective ground state electron densities. It is in the nature of self-consistent field (SCF) iterations that these initially make large jumps, resulting in an uneven training distribution that is mostly concentrated in a spike around the ground state density.
We instead modify the potential in the above equation to read $V_{\mathrm{eff}}[\{\phi_{i}^{t-1}\}]+\Delta^{t}$, where $\Delta^{t}\colon\mathbb{R}^{3}\to\mathbb{R}$ are randomly sampled perturbations. This approach results in more varied training labels (Fig. 3A and Fig. 3B which in turn afford the training of much better-behaved functionals, see Fig. 2 and ablation experiments in the supplementary text.
@@
The question of which target to train against is a highly interesting one. According to Hohenberg and Kohn [1], the total electronic energy can be decomposed into a universal (independent of the external potential) functional $F[\rho]$ and a classical electrostatic interaction between the electron density and the external potential representing the atomic nuclei: $E[\rho]=F[\rho]+\int\mathrm{d}^{3}\mathbf{r}\,v_{\mathrm{ext}}(\mathbf{r})\rho(\mathbf{r})$. Following Levy and Lieb [20], it is customary to further decompose the universal functional into a “non-interacting kinetic energy” $T_{\mathrm{S}}$, a classical electrostatic electron-electron or “Hartree” interaction $E_{\mathrm{H}}$ and an “exchange-correlation” term $E_{\mathrm{XC}}$, which captures all quantum effects not accounted for elsewhere. The immense practical success of Kohn-Sham DFT is owing to the fact that reasonably good approximations such as [22, 23, 24] to the true $E_{\mathrm{XC}}$ have been found, with active efforts underway to further improve upon these using machine learning [25].
At first sight, learning a density functional form of just $T_{\mathrm{S}}$, the raison d’être of Kohn-Sham DFT, seems natural in the spirit of reductionism. As an example, the model from [13] evaluated in Fig. 1A was trained on the difference of this non-interacting kinetic energy and the classical APBEK functional [9] in the fashion of “delta learning.”
@@
Finally, to obtain well-formed energy surfaces, it helps to use not only energies but also their functional derivatives as training targets. In the finite basis, the functional derivatives are gradients $\nabla_{\mathbf{p}}E$. Obtaining these is not trivial, and details can be found in materials and methods, see section A.1.
@@
A compact representation of electron density $\rho$ can be obtained in terms of a linear combination of atom-centered basis functions $\{\omega_{\mu}\}$ [14],
@@
We use an even-tempered Gaussian basis $\{\omega_{\mu}\}$ [26]. While this representation does not guarantee positivity, regions of unphysical negative densities are no failure case that we encounter in practice, see fig. S.2.
@@
Having experimented extensively with either class of architecture (e.g. [31, 32, 33]), we obtain broadly comparable results with representatives of both; but we currently find the best cost/performance tradeoff with a Graphormer [33, 34] type architecture. The latter profits from a self-attention mechanism, but is limited by the fact that it can only send scalar messages between nodes. In response, we invoke the formalism recently proposed in [19] to generalize the architecture to allow sending tensorial messages between nodes.
@@
The atomic basis functions $\{\omega_{\mu}\}$ overlap and so the density coefficients are not independent. We follow [13] in first transforming the coefficients into an orthonormal basis by means of a global “natural reparametrization” and then subjecting coefficients and energy gradients to dimension-wise rescaling and atomic reference modules, see [13] for details.
@@
Importantly, for larger molecules, we use a distance cutoff in the definition of atomic adjacency, making for a message passing mechanism that scales gracefully with system size. For detailed model specifications, see materials and methods.
Orbital-free DFT: A new status quo
Building on a compact representation of the electron density (Eq. 2, [14]), state-of-the-art machine learning architectures [33, 19], automated differentiation [35], efficient training data generation and reparametrization [13] and a strategy to create more balanced training samples [12], orbital-free DFT is finally starting to fulfill the promise that was implicit in the Hohenberg-Kohn theorems.
@@
We concentrate here on organic molecules with a mass of up to a few hundred Dalton. While they span only a tiny corner of the entire chemical space, this family is already estimated to comprise anywhere between $10^{24}$ and $10^{60}$ members [36, 37]. Needless to say, this number grows further when taking larger biopolymers such as peptides or nucleotide chains into account.
@@
The QM9 database [17], while restricted to stable molecules and relaxed geometries, is already quite diverse when considering only small organic molecules, see Fig. 4B. Orbital-free DFT as implemented by the STRUCTURES25 functional now affords density optimization which fully converges for each and every of the ca. 13k QM9 test molecules; with the resulting densities deviating from Kohn-Sham ground truth by around 0.46 electrons when integrated over all space. The biggest contribution to this deviation comes not from the machine learning model, but stems from the intrinsic error of density fitting, the process of expressing a Kohn-Sham density in the basis $\{\omega_{\mu}\}$ used to expand the orbital-free density in Eq. 2. The mean absolute energy errors of 0.64 mHa are well below the barrier of 1.6 mHa, a widely accepted definition of “chemical accuracy.”
@@
Reaching Kohn-Sham accuracy and fully convergent density optimization on the full chemistry embodied by the QM9 dataset is already most auspicious for orbital-free DFT. Yet, for future applications, reliable extrapolation to larger systems is essential: It is here that standard Kohn-Sham DFT becomes intractable. Following [13], we evaluate extrapolation accuracy on the QMugs database [38] comprising substantially larger druglike molecules. To this end, we train on smaller molecules with a mass up to around 200 Da, and evaluate on a subset of 850 test molecules, with a mass up to around 1400 Da. On these, the STRUCTURES25 functional achieves fully convergent density optimization on all molecules, shown in Fig. 4. Remarkably, even though the STRUCTURES25 functional uses message passing only across local neighborhoods up to six Bohr in radius, the mean absolute error per atom is larger, but does not grow with the size of the molecule, across the QMugs database (Fig. 4A). The three largest energy errors were all associated with the trifluoromethoxy group, a moiety which turned out to be represented only in terms of a single molecule in the training data. Larger molecules in the QMugs database highlight the improved scaling of machine-learned OF-DFT in comparison to the Kohn-Sham reference calculations: Minutes on a single Nvidia A100 GPU vs. hours on 10 CPU cores. Our quantitative results are summarized in Table 1.
Table 1: Orbital-free DFT finds ground state energies with chemical accuracy and meaningful densities on QM9 and QMugs. “Local” indicates whether the model guarantees a finite field of view. $|\Delta E|$: Mean absolute total energy error. $|\Delta E|/N$: Mean absolute energy error per atom. $\|\Delta\rho\|_{2}$: Mean $L_{2}$ norm of the density difference. $\|\Delta\rho\|_{2}/N_{e}$: Mean $L_{2}$ norm of the density difference, normalized by the number of electrons. Runtime is the average time for density optimization for a single molecule on an Nvidia Quadro RTX 6000 GPU for QM9 and an Nvidia A100 GPU for QMUGS.
Dataset Functional Target Local $|\Delta E|$ $|\Delta E|/N$ $\|\Delta\rho\|_{2}$ $\|\Delta\rho\|_{2}/N_{e}$ Runtime
(mHa) (mHa) ($\text{10}^{-\text{4}}$) (s)
QM9 STRUCTURES25 $E_{\mathrm{TXC}}$$\times$ 0.64 0.038 0.014 2.1 013
M-OFDFT $T_{\mathrm{S}}-\mathrm{APBEK}$$\times$ 1.37 0.088 0.027 4.2 183
QMugs STRUCTURES25 $E_{\mathrm{TXC}}$checkmark 22 0.21 0.068 1.6 040
M-OFDFT $E_{\mathrm{TXC}}$$\times$ 18 0.17 0.070 1.8 319
@@
What next for orbital-free DFT?
While immediate and worthwhile objectives are plentiful (including precise predictions on non-equilibrium geometries, in materials and open-shell systems, learning from higher-accuracy ground truth) we here concentrate on the scaling to huge systems.
Efficient scaling is also the subject of linear-scaling KS-DFT [39, 40, 41, 42]. To be meaningful, OF-DFT will not just need quasi-linear scaling, but also a small prefactor. Let us recall the computational complexity bottlenecks of OF-DFT. Initial density guesses such as MINAO [43, 44] pose a bottleneck as they come with cubic complexity in the number of atoms $N$. The robustness of our model regarding initialization allows us to start from our simple dSAD guess (see materials and methods) which scales linearly. Natural reparametrization still has $\mathcal{O}(N^{3})$ complexity and finding cheaper alternatives has top priority. Also, message passing on completely connected graphs comes with $\mathcal{O}(N^{2})$ complexity. Our local model operating on a radius graph reduces this to $\mathcal{O}(N)$. The cost of the Hartree term scales quadratically with the number of basis functions when the representation in Eq. 2 is evoked. Approximations such as the fast multipole method can bring this cost down to linearithmic scaling.
We expect future machine-learned density functionals to work well in those regimes in which Kohn-Sham DFT does. The viable area should be significantly larger than the kind of chemistries typified by the QM9 and QMugs databases.
Outside orbital-free DFT, direct property prediction using “foundation models” is progressing with great strides [45]. When striving to make valid predictions across a broad range of chemical space, a key question is which approach will prevail: A transductive approach, as in foundation models for property prediction which directly map from molecular geometry to observable; or an inductive ansatz, as in orbital-free DFT, where a universal functional is learned and variational optimization yields the electron density and its energy, from which observables can be derived. We believe the inductive approach will generalize better, but are eager to learn whether this intuition stands the test of time.
Overall, orbital-free DFT is now on the cusp of becoming practically useful in the molecular realm, thanks to the present contributions building on excellent prior work including [46, 33, 12, 13, 19].

### v2

@@ Abstract
Hohenberg and Kohn have proven that the electronic energy and the one-particle electron density can, in principle, be obtained by minimizing an energy functional with respect to the density. While decades of theoretical work have produced increasingly faithful approximations to this elusive exact energy functional, their accuracy is still insufficient for many applications, making it reasonable to try and learn it empirically. Using rotationally equivariant atomistic machine learning, we obtain for the first time a density functional that, when applied to the organic molecules in QM9, yields energies with chemical accuracy relative to the Kohn-Sham reference while also converging to meaningful electron densities. Augmenting the training data with densities obtained from perturbed potentials proved key to these advances. This work demonstrates that machine learning can play a crucial role in narrowing the gap between theory and the practical realization of Hohenberg and Kohn’s vision, paving the way for more efficient calculations in large molecular systems.
@@
Sadly, the proof of existence is non-constructive. That is, we know that a universal energy functional $F[\rho]$ exists; but its exact form remains unknown except for simple special cases [2, 3, 4] that do not cover most chemistries of general interest.
@@
The lure of a simple but complete description of molecular ground states in terms of their electron density alone has motivated an intense search for approximations with relatively few parameters [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. While remarkably successful considering the complex nature of the kinetic energy, these are not yet sufficiently accurate for the quantitative prediction of outcomes in typical laboratory chemistry.
This motivates the search for machine learned functionals whose significantly larger number of parameters promises greater expressivity [18, 19, 20]: While promising, initial forays [21, 22, 23, 24] were still limited regarding accuracy and/or transferability of the learned functionals. The equivariant KineticNet architecture [25] was the first to demonstrate chemical accuracy of a single learned functional across input densities and geometries of tiny molecules. This work was superseded by the milestone M-OFDFT pipeline [26]. Its representation of the density in terms of a linear combination of atomic basis functions [27, 28] is much more compact, compared to a grid. It first predicted molecular energies across the QM9 dataset [29, 30] of diverse organic molecules with up to nine second-period atoms with chemical accuracy with respect to the PBE/6-31G(2df,p) Kohn-Sham reference, a remarkable feat. Perhaps the biggest success of that work was its ability to extrapolate to larger molecules than trained on. One drawback was the model being fully non-local, i.e., each atom needs to exchange information with all others. While not a problem for molecules up to a few hundred atoms, this eventually becomes prohibitive for larger systems. Its foremost limitation, however, was that the learned functional did not afford variational density optimization: Following the energy gradient led to unphysical densities and energies, creating the need to engineer a method that picks an intermediate solution as the final prediction post hoc.
Figure 1: The STRUCTURES25 pipeline enables converging orbital-free electron density optimization. (A) Radar plot of total energy error, $L_{2}$ density error and percentage of convergence failures evaluated on QM9. STRUCTURES25 was trained on our perturbed QM9 dataset. Both energy and density error are w.r.t. PBE/6-31G(2df,p) ground state labels in the orbital-free density basis (Eq. 2). (B) The STRUCTURES25 pipeline takes a molecular graph $\mathcal{M}$ and a density represented by coefficients $\mathbf{p}$ of atom-centered basis functions $\{\omega_{\mu}\}$ as input and predicts the target energy, $E_{\mathrm{TXC}}$. The gradient of the energy is obtained by automatic differentiation and used to iteratively find the ground state in density optimization.
@@
Ensuring that the learned energy functional has a true minimum at the correct ground state electron density (or very nearby) is of paramount practical importance (Fig. 2A). Indeed, it enables convergent variational density optimization to meaningful densities (Fig. 2B). If no such minimum were present, following the gradient on such a misleading surface would result in unphysical densities, e.g., allowing electrons to collapse into a nucleus. Previously, such malformed energy surfaces required elaborate procedures to salvage a prediction from a diverging density optimization trajectory [26]. In addition, an estimated ground state density with vanishing gradients $\nabla_{\mathbf{p}}E=0$ is vital for the calculation of valid nuclear gradients, which are required for robust geometry optimization but especially for faithful molecular dynamics.
@@
yields a consistent tuple of potential $V_{\mathrm{eff}}$, orbitals $\phi_{i}$ and associated energies $\varepsilon_{i}$ from which a training sample (density coefficients $\mathbf{p}$, energy $E$, gradient $\nabla_{\mathbf{p}}E$) can be obtained. The bottom of Fig. 3C characterizes the resulting high-dimensional training data in terms of a single axis, the energy difference between training and respective ground state electron densities. It is in the nature of self-consistent field (SCF) iterations that these quickly converge to the vicinity of the ground state, resulting in an uneven training distribution that is mostly concentrated in a spike around the ground state density.
We instead modify the potential in the above equation to read $V_{\mathrm{eff}}[\{\phi_{i}^{t-1}\}]+\Delta^{t}$, where $\Delta^{t}\colon\mathbb{R}^{3}\to\mathbb{R}$ are randomly sampled perturbations. This approach results in more varied training labels (Figs. 3A and 3B) which in turn afford the training of much better-behaved functionals, see Fig. 2 and ablation experiments in supporting text S.10.
@@
The question of which target to train against is a highly interesting one. According to Hohenberg and Kohn [1], the total electronic energy can be decomposed into a universal (independent of the external potential) functional $F[\rho]$ and a classical electrostatic interaction between the electron density and the external potential representing the atomic nuclei: $E[\rho]=F[\rho]+\int\mathrm{d}^{3}\mathbf{r}\,v_{\mathrm{ext}}(\mathbf{r})\rho(\mathbf{r})$. Following Levy and Lieb [32], it is customary to further decompose the universal functional into a “non-interacting kinetic energy” $T_{\mathrm{S}}$, a classical electrostatic electron-electron or “Hartree” interaction $E_{\mathrm{H}}$ and an “exchange-correlation” term $E_{\mathrm{XC}}$, which captures all quantum effects not accounted for elsewhere. The immense practical success of Kohn-Sham DFT is owing to the fact that reasonably good approximations to the true $E_{\mathrm{XC}}$ have been found [34, 35, 36], with active efforts [37] underway to further improve upon these using machine learning.
At first sight, learning a density functional form of just $T_{\mathrm{S}}$, the raison d’être of Kohn-Sham DFT, seems natural in the spirit of reductionism. As an example, the model from M-OFDFT [26] evaluated in Fig. 1A was trained on the difference of this non-interacting kinetic energy and the classical APBEK functional [12] in the fashion of “delta learning.”
@@
Finally, to obtain well-formed energy surfaces, it helps to use not only energies but also their functional derivatives as training targets. In the finite basis, the functional derivatives are gradients $\nabla_{\mathbf{p}}E$. Obtaining these is not trivial, and details can be found in the supporting information, section S.1.
@@
A compact representation of electron density $\rho$ can be obtained in terms of a linear combination [27, 28] of atom-centered basis functions $\{\omega_{\mu}\}$,
@@
We use an even-tempered Gaussian basis [38] $\{\omega_{\mu}\}$. While this representation does not guarantee positivity, regions of unphysical negative densities are no failure case that we encounter in practice, see Fig. S.5 in the supporting information.
@@
Having experimented extensively with either class of architecture [43, 44, 45], we obtain broadly comparable results with representatives of both; but we currently find the best trade-off between cost and performance with a Graphormer [45, 46] type architecture. The latter profits from a self-attention mechanism, but is limited by the fact that it can only send scalar messages between nodes. In response, by invoking a recently proposed formalism [31], we generalize the architecture to allow sending tensorial messages between nodes.
@@
The atomic basis functions $\{\omega_{\mu}\}$ overlap and so the density coefficients are not independent. We follow M-OFDFT [26] in first transforming the coefficients into an orthonormal basis by means of a global “natural reparametrization” and then subjecting coefficients and energy gradients to dimension-wise rescaling and atomic reference modules, see [26] for details.
@@
Importantly, for larger molecules, we use a distance cutoff in the definition of atomic adjacency, making for a message passing mechanism that scales gracefully with system size. For detailed model specifications, see the supporting text.
Evaluating STRUCTURES25
@@
We concentrate here on equilibrium-geometry organic molecules with a mass of up to a few hundred Dalton. While they span only a tiny corner of the entire chemical space, this family is already estimated to comprise anywhere between $10^{24}$ and $10^{60}$ members [47, 48]. Needless to say, this number grows further when taking larger biopolymers such as peptides or nucleotide chains into account.
@@
The QM9 database [30], while restricted to stable molecules and relaxed geometries, is already quite diverse when considering only small organic molecules, see Fig. 4B. Orbital-free DFT as implemented by the STRUCTURES25 functional now affords density optimization which fully converges for each and every of the ca. 13k QM9 test molecules; with the resulting densities deviating from Kohn-Sham ground truth by $5.86\cdot 10^{-4}$ in the metric of the squared error integrated over space, normalized by the number of electrons. The biggest contribution to this deviation comes not from the machine learning model, but stems from the intrinsic error of density fitting (see Fig. S.2 in the supporting information), the process of expressing a Kohn-Sham density in the specific basis $\{\omega_{\mu}\}$ used to expand the orbital-free density in Eq. 2. Mean absolute energy errors of 0.64 mHa relative to the PBE/6-31G(2df,p) KS-DFT reference are well below the barrier of 1.6 mHa, a widely accepted definition of “chemical accuracy”.
@@
Reaching Kohn-Sham accuracy and fully convergent density optimization on the full chemistry embodied by the QM9 dataset is already most auspicious for orbital-free DFT. Yet, for future applications, reliable extrapolation to larger systems is essential: It is here that standard Kohn-Sham DFT becomes intractable. Following prior work [26], we evaluate extrapolation accuracy on the QMugs database [49] comprising substantially larger druglike molecules. To this end, we train on smaller molecules with a mass up to around 200 Da, and evaluate on a subset of 850 test molecules, with a mass up to around 1400 Da. On these, the STRUCTURES25 functional achieves fully convergent density optimization on all molecules, shown in Fig. 4. Remarkably, even though the STRUCTURES25 functional uses message passing only across local neighborhoods up to six Bohr in radius, the mean absolute error per atom is larger, but does not grow with the size of the molecule, across the QMugs database (Fig. 4A). The three largest energy errors were all associated with the trifluoromethoxy group, a moiety which turned out to be represented only in terms of a single molecule in the training data. Larger molecules in the QMugs database highlight the improved scaling of machine-learned OF-DFT in comparison to the Kohn-Sham reference calculations, resulting in nearly an order of magnitude speedup on identical hardware (section S.11 in the supporting information). While density optimization requires many more iterations than a typical Kohn-Sham SCF calculation, this is more than offset by the relative simplicity of each single iteration which does not require matrix diagonalization.
Table 1: Orbital-free DFT finds ground state energies with chemical accuracy relative
to a PBE/6-31G(2df,p) Kohn-Sham reference, and meaningful densities on QM9 and QMugs.
Dataset Functional Locala $|\Delta E|$b $|\Delta E|/N$c $\|\Delta\rho\|_{2}$d $\|\Delta\rho\|_{2}/N_{e}$e Runtimef
(mHa) (mHa) ($\text{10}^{-\text{2}}$) ($\text{10}^{-\text{4}}$) (s)
QM9 Ours $\times$ $\mathbf{0.644\boldsymbol{\pm}0.006}$ $\mathbf{0.0380\boldsymbol{\pm}0.0003}$ $\mathbf{1.40\boldsymbol{\pm}0.02}$ $\mathbf{2.12\boldsymbol{\pm}0.03}$ 013
M-OFDFTg $\times$ 1.37 0.088 2.7 4.2 183
QMugs Ours $\checkmark$ $25\pm 2$ $0.231\pm 0.019$ $\mathbf{6.8\boldsymbol{\pm}0.2}$ $\mathbf{1.61\boldsymbol{\pm}0.04}$ 040
M-OFDFT $\times$ 18 0.17 7.0 1.8 319
  • a
    Indicates whether the model guarantees a finite field of view.
  • b, c
    Mean absolute energy error, total and per atom.
  • d, e
    Mean $L_{2}$ norm of the density difference, total and per electron.
  • f
    Average time for density optimization for a single molecule on an Nvidia Quadro RTX 6000 GPU for QM9 and an Nvidia A100 GPU for QMUGS.
  • g
    Trained on the $T_{\mathrm{S}}-\mathrm{APBEK}$ target (all other models predict $E_{\mathrm{TXC}}$).
@@
Orbital-free DFT: Status quo, and where next
Building on a compact representation of the electron density [27, 28] (Eq. 2), state-of-the-art machine learning architectures [45, 31], automated differentiation [50], efficient training data generation and reparametrization [26] and a strategy to create more balanced training samples [25], orbital-free DFT is now emerging as a viable framework for accurate calculations grounded in the Hohenberg-Kohn theorems.
Still, important limitations persist. The present functional, trained on organic equilibrium-geometry molecules only, does not yet generalize to high-energy geometries (a prerequisite for successful geometry optimization), or across the periodic table. Also, the poor predictions of the trifluoromethoxy group illustrate that with the present size of training data, sufficient examples of chemical moietys must be seen during training. It will be fascinating to see to what extent future models, trained on more extensive data, will “grok” chemistry and generalize across more of chemical space. Indeed, we expect future machine-learned density functionals to work well in those regimes in which Kohn-Sham DFT does. In particular, the viable area should be significantly larger than the kind of chemistries typified by the QM9 and QMugs databases.
For the following discussion, we single out the question of scaling, one principal reason to study OF-DFT in the first place. To become an alternative to linear-scaling KS-DFT [51, 52, 53, 54], OF-DFT will not just need quasi-linear scaling, but also a small prefactor. Let us recall the computational complexity bottlenecks of OF-DFT, beginning with the initialization. While many classical OF-DFT functionals are sufficiently robust to start from cheap initial density guesses, the previous state of the art, the machine-learned M-OFDFT [26] was dependent on MINAO [55, 56] which has cubic complexity in the number of atoms $N$. The robustness of STRUCTURES25 allows starting from our simple dSAD guess (see supporting text) which scales linearly. The next major step in the computational pipeline is natural reparametrization, which still has $\mathcal{O}(N^{3})$ complexity. Finding cheaper alternatives has top priority. Next, message passing on completely connected graphs comes with $\mathcal{O}(N^{2})$ complexity. Our local model operating on a radius graph reduces this to $\mathcal{O}(N)$. The cost of the Hartree term scales quadratically with the number of basis functions when the representation in Eq. 2 is evoked. Approximations such as the fast multipole method can bring this cost down to linearithmic scaling.
Outside orbital-free DFT, direct property prediction using “foundation models” is progressing with great strides[57, 58, 59, 60]. When striving to make valid predictions across a broad range of chemical space, a key question is which approach will prevail: A transductive approach, as in foundation models for property prediction which directly map from molecular geometry to observable; or an inductive ansatz, as in orbital-free DFT, where a universal functional is learned and variational optimization yields the electron density and its energy, from which observables can be derived. We believe the inductive approach will generalize better, but are eager to learn whether this intuition stands the test of time.
Overall, orbital-free DFT is now on the cusp of becoming practically useful in the molecular realm, due in part to the present contributions building on excellent prior work [61, 45, 25, 26, 31].